OpenStudy (anonymous):

Juan will rent a car for the weekend. He can choose one of two plans. The first plan has an initial fee of $49.98 and costs an additional $0.16 per mile driven. The second plan has an initial fee of $57.98 and costs an additional $0.14 per mile driven. How many miles would Juan need to drive for the two plans to cost the same?

OpenStudy (anonymous):

juan to know the answer

OpenStudy (igreen):

We have two equations. y = 0.16x + 49.98 and y = 0.14x + 57.98

OpenStudy (igreen):

It's a system of equations. y = 0.16x + 49.98 y = 0.14x + 57.98 Solve the system of equations using substitution. 0.16x + 49.98 = 0.14x + 57.98 Solve for 'x'

OpenStudy (anonymous):

x=400

OpenStudy (anonymous):

right? @iGreen

OpenStudy (igreen):

Yes, that's correct 0.02x + 49.98 = 57.98 0.02x = 8 x = 400

OpenStudy (anonymous):

yaay!!

OpenStudy (igreen):

So Juan needs to drive 400 miles for the two plans to be the same.